A certain radioactive isotope has a half life of aprox 1600 years. How many years to the nearest year would be required for a given amount of this isotope to decay to 60% of that amount?
Answer by stanbon(75887)

A certain radioactive isotope has a half life of aprox 1600 years. How many years to the nearest year would be required for a given amount of this isotope to decay to 60% of that amount?
0.60A = A (1/2)^[t/1600]
0.60=(1/2)^[t/1600}
log(0.6)=[t/1600]log0.5
t/1600= log.6/log.5= 0.73696559...
t=1179.14... yrs.
Cheers,
Stan H.